OpenStudy (anonymous):

Simplify the expression. (8 + 7a) + 4 A. 12 + 11a B. 19a C. 8 + 11a D. 12 + 7a

OpenStudy (apoorvk):

Not again?

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 8 + 7a + 4 }\) Operate on like terms!

OpenStudy (anonymous):

8 + 7a + 4 = 8 + 4 + 7a = 12 + 7a D.